# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ VIII Discussion and Conclusion
Although CP shaving has become a common practice, the interactive influence of agents on CP time remains largely unexplored. This paper takes a first step toward understanding the application of the game framework in the context of CP shaving. We propose a theoretical game framework and show its applicability for the CP shaving problem by analyzing the game structure and analytically deriving the NE for each game structure under a two-agent two-period setting. We prove the equilibrium points are globally uniformly asymptotically stable, provided that all agents’ demand shifting remains within the baseline limit. We also show that the gradient-based algorithm, with an updating rule that approximates the asymptotically stable process via finite differences, can efficiently compute the equilibrium points. Our analysis reveals that agents’ strategic interactions in the game framework achieve the same peak-shaving effectiveness as an equivalent centralized peak-shaving model. However, under quasiconcave and non-concave conditions, the CP game results in a higher PoA. Our findings also persist for a general setting with more than two agents. Our insights on peak shaving effectiveness have the potential to inform utilities and operators that are concerned with applying game models.
@@
By analyzing the PoA in quasiconcave and non-concave games within the two-agent setting, we show that PoA increases with the disparities between agents, as measured by their marginal shifting costs. This insight informs the design of CP shaving mechanisms that simultaneously achieve effectiveness and fairness by aligning marginal shifting costs - either by incentivizing non-flexible agents with higher comfort loss or regulating them to shift less demand. We show that PoA increases as the game transitions from concave to non-concave and then to quasiconcave, corresponding to greater agent flexibility. This suggests that higher flexibility amplifies system inefficiency. Combined with our multi-agent analysis, which shows that PoA is sensitive to agent flexibility in small systems but stabilizes in larger ones, we suggest that forming large systems with flexible agents while grouping inflexible agents into smaller systems can effectively reduce inefficiency.
@@
Our work has a few limitations. First of all, we formulate the demand-shifting penalty as a quadratic function. Future research is needed to examine more generic functions and even obtain empirical models from data-driven methods. Another limitation is that we only look at a two-period scheduling problem with simultaneous decisions. Additional research is needed to study the CP game under a multi-period or a sequential context, where in the latter, agents demand shifting decisions are made stage by stage given the non-anticipatory price and demand realization. This also motivates the last limitation, which is that we did not consider the incomplete information of the game model. In practice, agents’ payoff structure should be private information that can’t be observed by others, but agents may get others’ previous decisions and system peak period to update their beliefs gradually. We would expect contextual optimization to be a promising way to infer the private payoff function from other observed features and embed it into the stochastic formulation.
@@
We separate three conditions to analyze the continuity and concavity due to the indicator function.
(1) We first analyze the concave game conditions. To make the game concave, the CP period can’t be changed, i.e., there must be only one period active in the entire gaming process. In this case, the indicator function is eliminated, and the payoff function returns to a quadratic function as described in (3), which is concave and continuous. Because changing CP period requires the sum of all agents’ demand shifting greater than the system balance point $b$, and the greatest demand shifting is their critical point. Thus, the conditions is $b<-r_{\mathrm{x}}-r_{\mathrm{y}}$ when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$ and $b>r_{\mathrm{x}}+r_{\mathrm{y}}$ when $S_{\mathrm{b},1}<S_{\mathrm{b},2}$.
(2) We then prove the quasiconcave/discontinuous conditions. When the CP period changes in the gaming process, it is obvious that the indicator function makes the game discontinuous. We then write the payoff function of agent x with the switching point $c_{\mathrm{x}}=b-y$
$f_{\mathrm{x}}(x,y)$ $=-(X_{1}+x)I(x-c_{\mathrm{x}})-(X_{2}-x)I(c_{\mathrm{x}}-x)-\alpha_{\mathrm{x}}x^{2},$ (27)
According to the definition of the quasiconcave function, for all $x^{\prime},x^{\prime\prime}\in\mathcal{X}_{\mathrm{x}}$ and $\lambda\in[0,1]$, agent x’s payoff function $f_{\mathrm{x}}(x,y)$ should satisfy the following for all $y\in\mathcal{X}_{\mathrm{y}}$,
$f_{\mathrm{x}}(\lambda x^{\prime}+(1-\lambda)x^{\prime\prime},y)\geq\min\{f_{\mathrm{x}}(x^{\prime},y),f_{\mathrm{x}}(x^{\prime\prime},y)\},$ (28)
which is determined by the switching point $c_{\mathrm{x}}$ and the critical point $r_{\mathrm{x}}$. There are three cases.
i) Agent x can switch CP period before reaching critical points in both periods 1 and 2, i.e., the switching point is within the critical point $-r_{\mathrm{x}}\leq c_{\mathrm{x}}\leq r_{\mathrm{x}}$, and we have
@@
Graphically, the function is monotonically increasing before the switching point and decreasing after the switching point.
ii) Agent x can only switch CP period before reaching critical points in period 1, i.e., the switching point is greater than the critical point of period 2 $r_{\mathrm{x}}<c_{\mathrm{x}}$. Also, the function is l.s.c. as defined in Definition 1, i.e., $-(X_{2}-c_{\mathrm{x}})<-(X_{1}+c_{\mathrm{x}})$, and we have
$r_{\mathrm{x}}<b-y,\ \frac{X_{2}-X_{1}}{2}=b_{\mathrm{x}}<b-y,$ (30)
Graphically, the function is quadratic (concave) in the left part of the switching point, monotonically decreasing in the right part of the switching point, and the function is l.s.c. in the switching point.
iii) Agent x can only switch CP period before reaching critical points in period 2, i.e., the switching point is less than the critical point of period 1 $-r_{\mathrm{x}}>c_{\mathrm{x}}$. Also, the function is u.s.c. as defined in Definition 1, i.e., $-(X_{2}-c_{\mathrm{x}})<-(X_{1}+c_{\mathrm{x}})$, and we have
$-r_{\mathrm{x}}>b-y,\ \frac{X_{2}-X_{1}}{2}=b_{\mathrm{x}}>b-y,$ (31)
Graphically, the function increases monotonically in the left part of the switching point, is quadratic (concave) in the right part of the switching point, and is u.s.c. in the switching point.
Combining these three scenarios, we need both agents’ payoff functions to satisfy one of them for all other agents’ strategies to make the game quasiconcave. When agent x satisfy case i), according to the condition (29), agent y’s strategy needs to satisfy $y\in[b-r_{\mathrm{x}},b+r_{\mathrm{x}}]$. From the graph description, we know agent y’s strategy is $y=c_{\mathrm{y}}=b-x$, $y=-r_{\mathrm{y}}$, and $y=r_{\mathrm{y}}$ under case i), ii), and iii), respectively. In case i), if $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$, agent x can balance its demand in the two periods, thus the maximum $x=b_{\mathrm{x}}$, if $b_{\mathrm{x}}>r_{\mathrm{x}}$ or $b_{\mathrm{x}}<-r_{\mathrm{x}}$, the best strategy is either $r_{\mathrm{x}}$ or $-r_{\mathrm{x}}$. This means $x\in[-r_{\mathrm{x}},r_{\mathrm{x}}]$, and thus $y\in[b-r_{\mathrm{x}},b+r_{\mathrm{x}}]$. Because both agents take the best strategy at the switching point, the only solution is $x=b_{\mathrm{x}},y=b_{\mathrm{y}}$, and taking this into (29), we have the conditions
@@
Otherwise, suppose agent x satisfies case ii), according to the condition (30), agent y’s strategy needs to satisfy $y<\min\{b-r_{\mathrm{x}},b-b_{\mathrm{x}}\}$. We know agent y’s strategy is $y=-r_{\mathrm{y}}$ and $y=r_{\mathrm{y}}$ under cases ii) and iii), and $r_{\mathrm{y}}$ are independent to $r_{\mathrm{x}}$ and $b_{\mathrm{x}}$. Thus, agent y under cases ii) and iii) can’t guarantee the quasiconcave conditions. We then conclude that each agent’s payoff function is quasiconcave if (32) is true, which shows each agent is capable according to Definition 3.
(3) Other than the above two conditions, the game is non-concave and still discontinuous due to the indicator function. In this case, at least one agent is non-capable, and they can change the CP period together. We can directly write the conditions as the complementary set of concave and quasiconcave conditions with respect to $\mathbb{R}$, i.e., (6d). Also, note that if one agent’s payoff function is quasiconcave or non-concave, the payoff functions of the other agents must also be quasiconcave or non-concave, which excludes the conditions that one agent’s payoff function is concave and the other is quasiconcave. This finishes the proof of this proposition. ∎
@@
The basic idea of proving this Theorem is to analyze whether agents are capable, if they are not, whether they are upper non-capable or lower-non-capable, and different conditions corresponding to different NE. The rationale is that both agents’ best response is to balance their demand in two periods, but this is limited by their baseline conditions, which is similar to analyzing the relationship between their critical point and balance point.
@@
Note that the lower non-capable agent x satisfies $b_{\mathrm{x}}<-r_{\mathrm{x}}$, corresponding to scenario iii) as described in the proof of Proposition 4. Graphically, it shows the function increases monotonically in the left part of the switching point, and quadratic (concave) in the right part of the switching point, but not u.s.c. in the switching point. While the upper non-capable agent x is $b_{\mathrm{x}}>r_{\mathrm{x}}$, corresponding to scenario ii) as described in the proof of Proposition 4. Graphically, it shows the function is quadratic (concave) in the left part of the switching point, monotonically decreasing in the right part of the switching point, but not l.s.c. in the switching point.
@@
We then separate many scenarios according to baseline demand conditions and whether they are non-capable agents, upper non-capable, or lower non-capable, to analyze the NE and corresponding conditions.
(1) $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$, i.e., baseline CP period is 1.
@@
(1a) $X_{1}\geq X_{2},Y_{1}\geq Y_{2}$, i.e., both agents’ individual demand is higher in the baseline CP period. We have four scenarios determined by whether they are capable or not.
  • Both agents are capable agents, i.e., $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq 0,-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0$, which means they can balance their demand, and the NE for all agent is their balance point,
    $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$ (33a)
    and we have the conditions for each agent as
    $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq 0,-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0.$ (33b)
  • Agent x is non-capable, while agent y is capable, i.e., $b_{\mathrm{x}}<-r_{\mathrm{x}}\leq 0,-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0$. Due to $X_{1}\geq X_{2}$, agent x must be lower non-capable, and the NE for agent x is the critical point, i.e., $x^{*}=-r_{\mathrm{x}}$. Agent y can at least shift demand until the system demands balance in the two periods, and as agent x can’t reach the balance point, agent y can save more by shifting more demand away from the CP period, but needs to compare with the shifting penalty. Thus, the NE for agent y is
    $y^{*}=\max\{-r_{\mathrm{y}},y^{\prime}\},y^{\prime}=b_{\mathrm{y}}-(x^{*}-b_{\mathrm{x}})=b-x^{*}\leq 0.$ (34a)
    Then we compare the $y^{\prime}$ with its critical point $-r_{\mathrm{y}}$. If $y^{\prime}<-r_{\mathrm{y}}$, i.e., $b-x^{*}=b+r_{\mathrm{x}}<-r_{\mathrm{y}}$, which is equivalent to $-r_{\mathrm{x}}-r_{\mathrm{y}}>b$, agent y can’t change the CP period solely, then $y^{*}=-r_{\mathrm{y}}$. The corresponding conditions for agent x is $-r_{\mathrm{x}}>b_{\mathrm{x}}$, and for agent y is
    $-r_{\mathrm{y}}\leq b_{\mathrm{y}}<-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}.$ (34b)
    where $-r_{\mathrm{y}}\leq b_{\mathrm{y}}$ is because agent y is capable. Otherwise, when $-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b$, the best strategy for agent y is $y^{\prime}$, and we have $y^{*}=b+r_{\mathrm{x}}$. The corresponding condition for agent x is still $-r_{\mathrm{x}}>b_{\mathrm{x}}$, for agent y is
    $b_{\mathrm{y}}\geq-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}.$ (34c)
    Note that this includes two scenarios when changing agent x to agent y.

### v5

@@ VIII Discussion and Conclusion
Although CP shaving has become a common practice, the interactive influence of agents on CP time remains largely unexplored. This paper takes a first step toward understanding the application of the game framework in the context of CP shaving. We propose a theoretical game framework and show its applicability for the CP shaving problem by analyzing the game structure and analytically deriving the NE for each game structure under a two-agent two-period setting. We prove that the equilibrium points are globally uniformly asymptotically stable, provided that all agents’ demand shifting remains within the baseline limit. We also show that the gradient-based algorithm, with an updating rule that approximates the asymptotically stable process via finite differences, can efficiently compute the equilibrium points. Our analysis reveals that agents’ strategic interactions in the game framework achieve the same peak-shaving effectiveness as an equivalent centralized peak-shaving model. However, under quasiconcave and non-concave conditions, the CP shaving game results in efficiency losses. Our findings also persist for a general setting with more than two agents. These theoretical findings are further validated through a real-world case study based on ERCOT’s 4CP program, demonstrating the practical relevance of the proposed game model.
@@
Our insights on peak shaving effectiveness have the potential to inform utilities and operators that are concerned with applying game models. By analyzing the efficiency loss in quasiconcave and non-concave games within the two-agent setting, we show that efficiency decreases with the disparities between agents, as measured by their marginal shifting costs. This insight informs the design of CP shaving mechanisms that simultaneously achieve effectiveness and fairness by aligning marginal shifting costs - either by incentivizing non-flexible agents with higher comfort loss or regulating them to shift less demand. We show that efficiency decreases as the game transitions from concave to non-concave and then to quasiconcave, corresponding to greater agent flexibility. This suggests that higher flexibility amplifies system inefficiency, which contradicts the common perception that utilities and operators generally favor flexible agents for their ability to provide grid services. Combined with our multi-agent analysis, which shows that efficiency loss is sensitive to agent flexibility in small systems but stabilizes in larger ones, we suggest that forming large systems with flexible agents while grouping inflexible agents into smaller systems can effectively increase efficiency.
@@
Our work has a few limitations. First of all, we formulate the demand-shifting penalty as a quadratic function. Future research is needed to examine more generic functions and even obtain empirical models from data-driven methods. Another limitation is that we only look at a two-period scheduling problem with simultaneous decisions. Additional research is needed to study the CP shaving game under a multi-period or a sequential context, where in the latter, agents demand shifting decisions are made stage by stage given the non-anticipatory price and demand realization. This also motivates the last limitation, which is that we did not consider the incomplete information of the game model. In practice, agents’ payoff structure should be private information that can’t be observed by others, but agents may get others’ previous decisions and system peak period to update their beliefs gradually. We would expect contextual optimization to be a promising way to infer the private payoff function from other observed features and embed it into the stochastic formulation.
Appendix: Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior
@@
We separate two conditions to analyze the continuity and concavity due to the indicator function.
(1) We begin by analyzing the concave game conditions. For the game to be concave, the CP period must always align with the baseline CP period throughout the game. This allows us to eliminate the indicator function, reducing the payoff function to the quadratic form described in (3), which is concave and continuous. Since altering the CP period requires the total demand shifting of all agents to exceed the system balance point $b$, and the maximum demand shifting within one period is limited by their respective critical points, the concavity condition is given by $b>r_{\mathrm{x}}+r_{\mathrm{y}}$. Similarly, if baseline CP period is 1, the condition is given by $b<-r_{\mathrm{x}}-r_{\mathrm{y}}$.
(2) We now establish the conditions for quasiconcave/discontinuous games. When the CP period shifts during the gaming process, the presence of the indicator function clearly introduces discontinuity in the game. To analyze this, we express agent x’s payoff function in terms of the switching point $c_{\mathrm{x}}=b-y$
$f_{\mathrm{x}}(x,y)$ $=-(X_{1}+x)I(x,y)-(X_{2}-x)(1-I(x,y))-\alpha_{\mathrm{x}}x^{2},$ (27a)
$I(x,y)=\begin{cases}1&x-c_{\mathrm{x}}\geq 0\\ 0&x-c_{\mathrm{x}}<0\end{cases},$ (27b)
According to the definition of quasiconcavity, for all $x^{\prime},x^{\prime\prime}\in\mathcal{X}_{\mathrm{x}}$ and $\lambda\in[0,1]$, agent x’s payoff function $f_{\mathrm{x}}(x,y)$ should satisfy the following for all $y\in\mathcal{X}_{\mathrm{y}}$,
$f_{\mathrm{x}}(\lambda x^{\prime}+(1-\lambda)x^{\prime\prime},y)\geq\min\{f_{\mathrm{x}}(x^{\prime},y),f_{\mathrm{x}}(x^{\prime\prime},y)\}.$ (28)
This quasiconcavity property is influenced by the switching point $c_{\mathrm{x}}$ and the critical point $r_{\mathrm{x}}$, leading to three possible cases for further analysis. Note that although by our consumption, the demand shifting $x$ of agent x is non-negative, the demand shifting $y$ of agent y is unrestricted.
i) Agent x can switch the CP period regardless of whether it occurs in period 1 or 2, meaning its switching point does not exceed its critical point in either period, i.e., $-r_{\mathrm{x}}\leq c_{\mathrm{x}}\leq r_{\mathrm{x}}$. This leads to the following conditions:
@@
Graphically, the payoff function is monotonically increasing before the switching point and decreasing afterward, forming a peak at $c_{\mathrm{x}}$
ii) Agent x can switch the CP period only if it occurs in period 1, meaning its switching point exceeds its critical point in period 2 but not in period 1, i.e., $r_{\mathrm{x}}<c_{\mathrm{x}}$. Additionally, the function satisfies the l.s.c. condition, as defined in Definition 1, i.e., $-(X_{2}-c_{\mathrm{x}})\leq-(X_{1}+c_{\mathrm{x}})$. This implies that
$r_{\mathrm{x}}<b-y,\ \frac{X_{2}-X_{1}}{2}=b_{\mathrm{x}}\leq b-y,$ (30)
Graphically, the function exhibits quadratic concavity to the left of the switching point, monotonic decrease to the right, and l.s.c. at the switching point.
iii) Agent x can switch CP period only if it occurs in period 2, meaning its switching point exceeds its critical point in period 1 but not in period 2, i.e., $-r_{\mathrm{x}}>c_{\mathrm{x}}$. Also, the function satisfies the u.s.c. conditions as defined in Definition 1, i.e., $-(X_{2}-c_{\mathrm{x}})\leq-(X_{1}+c_{\mathrm{x}})$. This implies that
$-r_{\mathrm{x}}>b-y,\ \frac{X_{2}-X_{1}}{2}=b_{\mathrm{x}}\geq b-y,$ (31)
Graphically, the function exhibits quadratic concavity to the right of the switching point, monotonic decrease to the left, and u.s.c. at the switching point.
By combining the three cases, we require that at least one of them holds for both agents’ payoff functions across all possible strategies of the other agent to ensure the game remains quasiconcave. From the graphical representation, we observe that if an agent’s payoff function satisfies case i), the optimal strategy is at its switching point $x=c_{\mathrm{x}},y=c_{\mathrm{y}}$. If an agent’s payoff function satisfies case ii) or iii), the optimal strategy is at its critical point, specifically $x=r_{\mathrm{x}},y=r_{\mathrm{y}}$ for case ii) and $x=-r_{\mathrm{x}},y=-r_{\mathrm{y}}$ for case iii). Now, suppose agent x’s payoff function satisfies case i). According to condition (29), agent y’s strategy needs to satisfy both $y\geq b-r_{\mathrm{x}}$ and $y\leq b+r_{\mathrm{x}}$. If agent y’s payoff function satisfies case ii) or iii), we obtain either $r_{\mathrm{x}}+r_{\mathrm{y}}\geq b$ or $-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b$, which correspond to the concave game conditions as previously described, except when equality holds, i.e., $r_{\mathrm{x}}+r_{\mathrm{y}}=b$ or $-r_{\mathrm{x}}-r_{\mathrm{y}}=b$. These equality conditions imply that their critical points equal to their balance points, meaning $b_{\mathrm{x}}=r_{\mathrm{x}}$ and $b_{\mathrm{y}}=r_{\mathrm{y}}$. Otherwise, for the game to be quasiconcave, both agents’ payoff functions must satisfy case (i). This implies that both agents can fully balance their demand over the two periods, and their maximum demand shifting are given by $x=b_{\mathrm{x}},y=b_{\mathrm{y}}$. Substituting these into condition (29), we derive the conditions that ensure quasiconcavity.
@@
Otherwise, suppose agent x’s payoff function satisfies case ii), according to the condition (30), agent y’s best strategy needs to satisfy $y<b-r_{\mathrm{x}},y\leq b-b_{\mathrm{x}}$. We know agent y’s best strategy is $y=-r_{\mathrm{y}}$ or $y=r_{\mathrm{y}}$ when its payoff function satisfies cases ii) or iii), respectively. As $r_{\mathrm{y}}$ are independent to $r_{\mathrm{x}}$ and $b_{\mathrm{x}}$, agent y’s payoff function satisfies cases ii) or iii) can’t guarantee the quasiconcave conditions. Thus, we conclude that each agent’s payoff function is quasiconcave if (32) is true, which also shows each agent is capable according to Definition 3. Given our baseline assumption, agent x’s condition can further simplified to $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$.
(3) Other than the above two conditions, the game is non-concave and still discontinuous due to the indicator function. In this case, at least one agent is non-capable, and they can change the CP period together. We can directly write the conditions as the complementary set of concave and quasiconcave conditions with respect to $\mathbb{R}$, i.e., (6c). Also, note that if one agent’s payoff function is quasiconcave or non-concave, the payoff functions of the other agents must also be quasiconcave or non-concave, which excludes the conditions that one agent’s payoff function is concave and the other is quasiconcave. This finishes the proof of this proposition. ∎
@@
Overview of the proof: We begin by deriving a Lemma to establish the NE for the quasiconcave CP shaving game $G$, providing insight into the best response rationale. We then proceed to prove the theorem by analyzing the shifting capabilities of both agents.
Lemma 1.
NE under quasiconcave two-agent setting. Under the quasiconcave two-agent, two-period setting as described in Proposition 1, the NE of the CP shaving game, as defined in Definition 1, is given by $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{x}}$ and satisfies the condition $S_{1}(x^{*},y^{*})=S_{2}(x^{*},y^{*})$.
@@
Since we are dealing with the quasiconcave condition, we assume $x$ is non-negative, while $y$ remains unrestricted. Thus, we conduct a comprehensive analysis without imposing specific baseline demand assumptions. From the best response perspective, for a given instance, suppose the CP period is 1, meaning $S_{1}=S_{2}+\delta$, where $\delta$ is an infinitesimal number. If agent x benefits by reducing $x$, shifting the CP period to 2, this implies that agent x’s demand in period 1 is greater than in period 2, i.e., $X_{1}+x\geq X_{2}-x$. Now, considering agent y, if its demand in period 1 is lower than in period 2 ($Y_{1}+y<Y_{2}-y$), the CP period shift from 1 to 2 harms agent y, leading them to increase $y$ in an attempt to push the CP period back to 1. Conversely, if $Y_{1}+y\geq Y_{2}-y$, shifting the CP period to 2 benefits agent y as it can shift $y$ away from period 2, meaning its best response remains to increase $y$. Thus, regardless of agent y’s demand distribution, its best response is always adversarial to agent x.
This demonstrates that both agents are fully competitive—if agent x benefits from changing their strategy, that same strategy directly harms agent y. This mirrors a zero-sum game setting, where one agent’s gain is the other’s loss. Given this competitive structure, we can apply the following min-max formulation to analyze their best responses.
where $y^{*}$ is the best strategy of agent y following the same structure, and the first and second case corresponds to CP periods 1 and 2. By applying the first-order optimality condition, the best strategy of $x$ is
The best strategy for agent y follows the same structure
According to (6b), we know $-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$, $-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}$, then the (33b) and (33c) can be simplified to $y^{*}=\max\{b-x^{*},b_{\mathrm{y}}\}$ and $x^{*}=\max\{b-y^{*},b_{\mathrm{x}}\}$, which shows the only mutual best strategy are $x^{*}=b_{\mathrm{x}},\ y^{*}=b_{\mathrm{y}}$ (the NE), and the conditions are
$X_{2}+Y_{2}-x^{*}-y^{*}=S_{2}=X_{1}+Y_{1}+x^{*}+y^{*}=S_{1}.$ (34)
This shows the NE will always be obtained when $S_{1}=S_{2}$ under the quasiconcave condition. Essentially, during the game, periods 1 and 2 are active interactively and finally converges at the connecting point of both periods, i.e., $S_{1}=S_{2}$. This means there is a unique equilibrium point in the entire strategy set $\mathcal{X}$. Indeed, all agents will minimize the payment associated with $S_{1},S_{2}$ imbalance as any imbalance results in a significant CP charge change caused by the opponent’s strategy. ∎
This lemma establishes the NE under the quasiconcave game condition and highlights that agents’ best response is to adopt adversarial strategies when their individual peak periods differ to reduce their CP charge. In contrast, the concave game is straightforward to analyze, as both agents maximize quadratic independent payoff functions, making their best strategies simply their critical points of the CP period. Building on this foundation, we now present the proof of the theorem.
Proof of Theorem 1.
@@
The basic idea behind proving this theorem is to analyze whether agents are capable of fully balancing their demand. If they are not, we further determine whether they are upper non-capable or lower non-capable, as each condition corresponds to a different NE. The key rationale is that both agents’ best response is to shift demand to reduce CP charge, but their shifting capability imposes a constraint. Thus, the key aspect of the proof lies in analyzing the relationship between their critical point and balance point.
@@
Given our assumption that $S_{\mathrm{b},1}<S_{\mathrm{b},2}$ and $X_{1}<X_{2}$, if agent x is non-capable, it must be upper non-capable, satisfying $b_{\mathrm{x}}>r_{\mathrm{x}}$. This corresponds to scenario ii) as described in Proposition 1. Graphically, this means that agent x’s payoff function is quadratic (concave) in the left part of the switching point, monotonically decreasing in the right part, but does not satisfy l.s.c. at the switching point. Meanwhile, agent y remains unrestricted in its decision-making.
@@
We then separate many scenarios according to whether they are non-capable agents, upper non-capable, or lower non-capable, to analyze the NE and corresponding conditions.
@@
(1) $X_{1}<X_{2},Y_{1}<Y_{2}$, meaning both agents have higher individual demand in the baseline CP period, which is period 2 by assumption. The analysis is divided into four scenarios, depending on whether each agent is capable or non-capable.
  • Both agents are capable agents, i.e., $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},0\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}$, the game is quasiconcave, meaning both agents can fully balance their demand over two periods. By Lemma 1, the NE is $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$.
  • Agent x is non-capable while agent y is capable, i.e., $b_{\mathrm{x}}>r_{\mathrm{x}}>0,0\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}$, agent x must be upper non-capable since $X_{1}<X_{2}$, and its best strategy is to shift demand up to its critical point in period 2, so, $x^{*}=r_{\mathrm{x}}$. Since agent y is capable, it can further shift demand to help balance the system. Essentially, agent y can reduce costs by shifting demand away from CP period 2 but must do so within its shifting capability. The optimal strategy for agent y depends on whether it can fully balance the system demand given agent x’s strategy. Thus, the best strategy for agent y is $y^{*}=\min\{r_{\mathrm{y}},b-x^{*}\}$. This distinction determines whether the game remains concave or becomes non-concave, leading to two scenarios. To determine the scenario, we compare the two terms. i) If $b-x^{*}=b-r_{\mathrm{x}}>r_{\mathrm{y}}$, meaning agent y cannot shift enough to change the CP period given agent x’s strategy, then the game remains concave, satisfying $r_{\mathrm{x}}+r_{\mathrm{y}}<b$. In this case, the best strategy is $y^{*}=r_{\mathrm{y}}$, with the conditions $r_{\mathrm{x}}<b_{\mathrm{x}},b_{\mathrm{y}}\leq r_{\mathrm{y}}$ and $r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{x}}<b_{\mathrm{y}}$ ii) Otherwise, when $r_{\mathrm{x}}+r_{\mathrm{y}}\geq b$, agent y can shift enough demand to change the CP period, making the game non-concave. In this scenario, agent y’s best strategy is $y^{*}=b-x^{*}=b-r_{\mathrm{x}}$. The corresponding condition is $r_{\mathrm{x}}<b_{\mathrm{x}},b_{\mathrm{y}}\leq r_{\mathrm{y}}$ and $b_{\mathrm{y}}\leq r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{x}}$.
  • Agent x is capable while agent y is non-capable, i.e., $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},b_{\mathrm{y}}>r_{\mathrm{y}}$. Similar to prior scenario, agent y can only be upper non-capable. Thus, by symmetry, we can interchange x and y to derive the NE and corresponding conditions. The equilibrium strategies are $y^{*}=r_{\mathrm{y}},x^{*}=r_{\mathrm{x}}$ under the condition of $b_{\mathrm{x}}\leq r_{\mathrm{x}},b_{\mathrm{y}}>r_{\mathrm{y}}$ and $r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}<b_{\mathrm{x}}\leq r_{\mathrm{x}}$, while $y^{*}=r_{\mathrm{y}},x^{*}=b-r_{\mathrm{y}}$ under the condition of $b_{\mathrm{x}}\leq r_{\mathrm{x}},b_{\mathrm{y}}>r_{\mathrm{y}}$ and $b_{\mathrm{x}}\leq r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}$;